# dApps: Enabling Real-Time AI-Based Open RAN Control

### v1

@@ 6.1 Use Case: Spectrum Sharing
Results in Table 2 show that the execution of the dApp (cases where dApp = Y and Incumbent = N) does not introduce any significant impact on throughput if compared to the case where the dApp is not enabled (cases where dApp = N and Incumbent = N). Specifically, in the Colosseum testbed, the average throughput measured without the dApp is 71.34 Mbps, while we measure an average throughput of 71.53 Mbps when the dApp is enabled. Similarly, in the Arena testbed, the throughput decreases slightly from 77.98 Mbps without the dApp to 76.37 Mbps with the dApp. Despite the average values seem to suggest a performance degradation, we notice that confidence intervals reported in Table 2 suggest that the throughput remains within the same order of magnitude, demonstrating that the dApp introduces minimal overhead in these scenarios that allows for performance parity.
@@ 6.2 Use Case: Sensing and Positioning
The above described ranging dApp framework is tested and validated with real-world experiments. The experimental setup consists of a single antenna gNB and a UE communicate over a line-of-sight channel in an indoor lab environment as shown in Fig. 11. The gNB and the UE rely on the OAI protocol stack and USRP B210 software-defined boards. Furthermore, the SC2430 NR signal conditioning module is used as an external RF front-end at the gNB [51]. The system operates at band n78 with a carrier frequency of $3.69$ GHz and a bandwidth of 40 MHz. The UL CIR extracted by the ranging dApp is shown in Fig. 11.
The channel gain and channel impulse response both depend on UE ’s position because path loss, shadowing, and multi-path are all phenomena that vary with distance and location of the UE. Closer UEs benefit from lower path loss and thus higher gain, while changes in nearby obstacles and reflections impact the channel impulse response by altering the number and timing of multi-path components. As a UE moves, both large-scale fading (due to environmental obstructions) and small-scale fading (due to the interference of different signal paths) continually vary the channel characteristics, causing the gain and impulse response to change accordingly. For details on ranging algorithms and prototype design, we refer the readers to [50].
Figure 11: Setup for the Sensing and Positioning dApp and UL CIR extracted by the dApp
7 Related Work and Comparison with Real-time RIC
The need for the introduction of real-time control loops in O-RAN has been advocated in recent works to control lower-layer functionalities of the gNB protocol stack [22, 52]. In our previous work [15], we introduced the concept of dApps to process data locally available at the gNB, and the need for sub-$10$ ms control loops. However, our work primarily focused on demonstrating the benefits of such tighter control loops for generic use cases rather than on proposing a complete reference architecture, as our work instead does.
The need for real-time control loops and inference in O-RAN has also been advocated by the broader community. For example, the concept of Real-time RICs and its applications have been proposed in [17, 18, 16]. Specifically, [17] proposes to embed CUs and DUs inside the real-time RIC, as well as to further disaggregate their functionalities, where each one (e.g., MAC scheduler, channel estimators, and beam shapers) is carried out by a dedicated application, namely zApp. [18] proposes TinyRIC, which runs close to the CU/DU, and so-called tApps hosted therein. The latter aid the CU/DU by performing functionalities spanning from the management of non-3GPP interfaces (e.g., O1 and E2), to energy management, to providing primitives for data collection. Similarly, [16] proposed EdgeRIC and its applications, $\mu$Apps, that interface with the RAN to make decisions at Transmission Time Interval (TTI) level. Although these approaches propose a viable way to introduce real-time control in the RAN, they do so by replicating functionalities typical of non- and Near-RT RICs, e.g., intelligence orchestration, in a new RT RIC platform, which inherently increases the network complexity, as well as resource utilization and energy consumption. Indeed, the main difference between the above-mentioned RT applications and dApps lies in the way intelligent applications are hosted and executed. A high-level architectural comparison of the two approaches is shown in Fig. 12, which highlights the placement of intelligent applications in the each of them.
(a) RT-RIC deployment
(b) dApp deployment
Figure 12: Comparison between RT RIC and dApp architectures. From [19].
The former require a dedicated RT RIC entity with an additional layer of abstraction of the network, while the latter can be executed at the CU/DU directly, minimize the impact on the E2 nodes and taking advantage of interfaces and functionalities already standardized by O-RAN.
Finally, [53] proposes Janus, a monitoring and control framework that can load and execute custom real-time “codelets” on the CU and DU. Even though this work—possibly the most similar to ours—extends O-RAN to real-time control, Janus’s codelets directly modify the protocol stack of CU/DU, instead interfacing with it. In addition, the current version of Janus only works with CPU-based stacks. Moreover, this approach also requires the definition and deployment of a dedicated controller on the RIC, instead of relying on standardized components for the management of real-time RAN functions, as we instead do in our dApp architecture.
@@ 8 Conclusions
In this paper, we introduced an extension of the O-RAN architecture focused on enabling (i) real-time control in the RAN, and (ii) the interaction of the O-RAN stack with the user-plane of the network. This is achieved through dApps, lightweight microservices co-located with DUs and CUs, and holistically managed through components such as the RICs and SMO. We first described the role that dApps have in enabling new use cases for RAN optimization, including inference based on I/Q samples or reference signals available at the physical layer. Based on the data and control requirements defined by these use cases, we discussed the architecture for the integration of dApps and RAN nodes. We also provided insights on the design and implementation of a set of APIs that RAN nodes need to expose to dApps, coordinated by an API broker called E3 agent. We then discussed the LCM for dApps, including steps for development, onboarding, and deployment of dApps, as well as procedures for interaction with the RICs and SMO.
@@
For the positioning use case, the dApp processes uplink reference signals to extract the uplink channel response and perform ranging for the UE. We profiled the performance of our dApp framework, demonstrating that real-time control loops under 1 ms are achievable in O-RAN. In our implementation, the average control loop duration is slightly less than 400 $\mu$s, on average.
@@ Acknowledgement
This article is based upon work partially supported by OUSD(R&E) through Army Research Laboratory Cooperative Agreement Number W911NF-24-2-0065. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. The work was also partially supported by the O-RAN ALLIANCE, the SERICS (PE00000014) 5GSec project under the MUR National Recovery and Resilience Plan funded by the European Union — NextGenerationEU, and by the U.S. National Science Foundation under grants CNS-1925601, CNS-2117814, and CNS-2112471.

### v2

@@ 7.1 Use Case: Spectrum Sharing
Results in Table 3 show that the execution of the dApp (cases where dApp = Y and Incumbent = N) does not introduce any significant impact on throughput if compared to the case where the dApp is not enabled (cases where dApp = N and Incumbent = N). Specifically, in the Colosseum testbed, the average throughput measured without the dApp is 71.34 Mbps, while we measure an average throughput of 71.53 Mbps when the dApp is enabled. Similarly, in the Arena testbed, the throughput decreases slightly from 77.98 Mbps without the dApp to 76.37 Mbps with the dApp. Although the average values may suggest performance degradation, the confidence intervals in Table 3 indicate that throughput takes values within the same confidence range, demonstrating that the dApp introduces minimal overhead and maintains performance parity.
@@ 7.2 Use Case: Sensing and Positioning
Figure 12: Setup for the Sensing and Positioning dApp and UL CIR extracted by the dApp.
@@
The above described ranging dApp framework is tested and validated with real-world experiments. The experimental setup consists of a single antenna gNB and a UE communicate over a line-of-sight channel. The gNB and the UE rely on on the OAI protocol stack and USRP B210 software-defined boards. Furthermore, the SC2430 NR signal conditioning module is used as an external RF front-end at the gNB [54]. The system operates at band n78 with a carrier frequency of $3.69$ GHz and a bandwidth of 40 MHz.
Performance Evaluation – The performance of the proposed ranging scheme is evaluated in terms of empirical CDF of the distance estimation error at an UL SNR of 20-20 dB. The CDF is obtained from 48,000 CIR measurements by fixing the position of the gNB and moving the UE in a straight line between 3 to 10 meters with a 1-meter increment. At every point, a total of 6,000 measurements are collected. The ranging dApp is used to extract these measurements as shown in Fig. 12. The CDF of the range estimation error with a super resolution Multiple Signal Classification (MUSIC) algorithm with M=20M{=}20 to $60$ number of CIR measurements is shown in Figure 13. For details on ranging algorithms and prototype design, we refer the readers to [53]. While the ranging dApp here is used to collect measurements during the experiment, which are then processed offline to obtain the range error CDF, it’s important to highlight that it can also be used for real-time CIR collection and distance evaluation. The measurement collection latency performance is similar to that described in Section 6.3.
@@ 8 Conclusions
In this paper, we introduced an extension of the O-RAN architecture focused on enabling (i) real-time control in the RAN, and (ii) the interaction of the O-RAN stack with the user-plane of the network. This is achieved through dApps, lightweight microservices co-located with DUs and CUs, and holistically managed through components such as the RICs and SMO. We first described the role that dApps have in enabling new use cases for RAN optimization, including inference based on I/Q samples, reference signals available at the physical layer, or as complex ISAC systems. Based on the data and control requirements defined by these use cases, we discussed the architecture for the integration of dApps and RAN nodes. We also provided insights on the design and implementation of a set of APIs that RAN nodes need to expose to dApps, coordinated by an API broker called E3 agent. We then discussed the LCM for dApps, including steps for development, onboarding, and deployment of dApps, as well as procedures for interaction with the RICs and SMO.
@@
[…]
@@ Acknowledgement
This article is based upon work partially supported by OUSD(R&E) through Army Research Laboratory Cooperative Agreement Number W911NF-24-2-0065. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. The work was also partially supported by SERICS (PE00000014) 5GSec project, CUP B53C22003990006, under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU, and by the U.S. National Science Foundation under grants CNS-1925601, CNS-2117814, and CNS-2112471.